Libtasn1

Abstract Syntax Notation One (ASN.1) library for the GNU system for version 4.1, 27 April 2014

Fabio Fiorina Simon Josefsson Nikos Mavrogiannopoulos (help-libtasn1@gnu.org)

This manual is for GNU Libtasn1 (version 4.1, 27 April 2014), which is a library for Abstract Syntax Notation One (ASN.1) and Distinguished Encoding Rules (DER) manipulation. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2001-2014 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 ASN.1 structure handling . . . . . . . . . . . . . . . . . . . . . . 2

2.1 ASN.1 syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Naming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Simple parsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.4 Library Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Future developments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Invoking asn1Parser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Invoking asn1Coding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Invoking asn1Decoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4 Function reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4.1 ASN.1 schema functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.2 ASN.1 field functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3 DER functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.4 Error handling functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.5 Auxilliary functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

Appendix A Copying Information . . . . . . . . . . . . . 22

A.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

Function and Data Index . . . . . . . . . . . . . . . . . . . . . . . . . 31

1 Introduction

This document describes the Libtasn1 library that provides Abstract Syntax Notation One (ASN.1, as specified by the X.680 ITU-T recommendation) parsing and structures management, and Distinguished Encoding Rules (DER, as per X.690) encoding and decoding functions.

The main features of this library are:

• On-line ASN.1 structure management that doesn’t require any C code file generation.

• Off-line ASN.1 structure management with C code file generation containing an array.

• Distinguished Encoding Rules (DER) encoding support.

• No limits for INTEGER and ENUMERATED values.

• It’s Free Software. Anybody can use, modify, and redistribute the library under the terms of the GNU Lesser General Public License version 2.1 or later. The command line tools, self-tests and build infrastructure are licensed under the GNU General Public License version 3.0 or later.

• Thread-safety. No global variables are used and multiple library handles and session handles may be used in parallel.

• Portability. The code should work on all Unix like operating systems, and Windows. The library itself should be portable to any C89 system, not even POSIX is required.

2 ASN.1 structure handling

2.1 ASN.1 syntax

The parser is case sensitive. The comments begin with -- and end either with another --, or at the end of the respective line, whichever comes first. The C-style /*, */ comments are not supported.

For an example of the syntax, check the pkix.asn file distributed with the library. ASN.1 definitions must follow the syntax below:

    definitions_name {<object definition>}
    DEFINITIONS <EXPLICIT or IMPLICIT> TAGS ::=
    BEGIN
    <type and constants definitions>
    END

The ::= token must be separate from other elements, so the following declaration is invalid:

      -- INCORRECT
      Version ::=INTEGER

The correct form is:

      Version ::= INTEGER

Here is the list of types that the parser can manage:

• INTEGER;
• ENUMERATED;
• BOOLEAN;
• OBJECT IDENTIFIER;
• NULL;
• BIT STRING;
• OCTET STRING;
• UTCTime;
• GeneralizedTime;
• GeneralString;
• NumericString;
• IA5String;
• TeletexString;
• PrintableString;
• UniversalString;
• BMPString;

• UTF8String;
• VisibleString;
• SEQUENCE;
• SEQUENCE OF;
• SET;
• SET OF;
• CHOICE;
• ANY;
• ANY DEFINED BY.

This version doesn’t handle the REAL type. It doesn’t support the AUTOMATIC TAGS option, and the EXPORT and IMPORT sections, either.

The SIZE constraints are allowed, but no check is done on them.

2.2 Naming

Consider this definition:

    Example { 1 2 3 4 }
    DEFINITIONS EXPLICIT TAGS ::=
    BEGIN
    Group ::= SEQUENCE {
       id   OBJECT IDENTIFIER,
       value  Value
    }
    Value ::= SEQUENCE {
       value1  INTEGER,
       value2  BOOLEAN
    }
    END

The notation to access the ‘Group’ type of the ‘Example’ definition above is ‘Example.Group’ (as a NUL-terminated string.) Such strings are used in the functions described below.

Others examples:

• field ‘id’ of the ‘Group’ type: ‘Example.Group.id’;
• field ‘value1’ of the ‘value’ field of the ‘Group’ type: ‘Example.Group.value.value1’.

Elements of structured types unnamed by the respective definition receive the names ?1, ?2, and so on.

The ?LAST name indicates the last element of a SET OF or SEQUENCE OF.

2.3 Simple parsing

For simple types like OCTET STRING the simple parsing functions listed below may be used instead.

• [asn1 decode simple der], page 20

• [asn1 encode simple der], page 15

2.4 Library Notes

The header file of this library is libtasn1.h.

The main type used in it is asn1_node, and it’s used to store the ASN.1 definitions and structures (instances).

The NULL constant can be used for the variable initialization. For example:

      asn1_node definitions = NULL;

Some functions require an errorDescription argument of type char *, pointing to a pre-allocated buffer of at least ASN1_MAX_ERROR_DESCRIPTION_SIZE bytes size (e.g., as in

‘char description[ASN1_MAX_ERROR_DESCRIPTION_SIZE];’).

ASN1_MAX_NAME_SIZE is the maximum number of characters allowed for an ASN.1 identifier.

2.5 Future developments

• Add functions for a C code file generation containing equivalent data structures (not a single array like now).

• The REAL type.

3 Utilities

3.1 Invoking asn1Parser

asn1Parser reads a single file with ASN.1 definitions and generates a file with an array to use with libtasn1 functions.

Usage:  asn1Parser [options] file
Options:
 -h : shows the help message.
 -v : shows version information and exit.
 -c : checks the syntax only.
 -o file : output file.
 -n name : array name.

3.2 Invoking asn1Coding

asn1Coding generates a DER encoding from a file with ASN.1 definitions and another one with assignments.

The file with assignments must have this syntax:

InstanceName  Asn1Definition
nameString  value
nameString  value
...

To specify the field of a CHOICE to be used, specify its name as a value to the CHOICE element itself. Use '' to denote the root element itself. (as in the example below.)

The output file is a binary file with the DER encoding.

Usage:  asn1Coding [options] file1 file2
 file1 : file with ASN1 definitions.
 file2 : file with assignments.
Options:
 -h : shows the help message.
 -v : shows version information and exit.
 -c : checks the syntax only.
 -o file : output file.

For example, consider an ASN.1 definitions file as follows:

PKIX1 { }
DEFINITIONS IMPLICIT TAGS ::=
BEGIN
Dss-Sig-Value ::= SEQUENCE {

     r       INTEGER,
     s       INTEGER
}
END

And a assignments file as follows:

dp PKIX1.Dss-Sig-Value
r 42
s 47

Running the command below will generate a assign.out file, containing the DER en-

coding of PKIX1.Dss-Sig-Value.
$ asn1Coding pkix.asn assign.asn1

If the root element is of the CHOICE type, the assignment file may be like (using the types defined in pkix.asn):

elt PKIX1Implicit88.GeneralName
''      dNSName
dNSName example.org

3.3 Invoking asn1Decoding

asn1Decoding generates an ASN.1 structure from a file with ASN.1 definitions and a binary file with a DER encoding.

Usage:  asn1Decoding [options] file1 file2 type
 file1 : file with ASN1 definitions.
 file2 : binary file with a DER encoding.
 type : ASN1 definition name.
Options:
 -h : shows the help message.
 -v : shows version information and exit.
 -o file : output file.

For example, after generating the assign.out file from the example section of the asn1Coding command above, the following invocation will decode the DER data.

$ asn1Decoding pkix.asn assign.out PKIX1.Dss-Sig-Value

4 Function reference

4.1 ASN.1 schema functions

asn1 parser2tree

[Function] int asn1_parser2tree (const char * file, asn1 node * definitions,
   char * error_desc)

file: specify the path and the name of file that contains ASN.1 declarations.

definitions: return the pointer to the structure created from "file" ASN.1 declarations.

error desc: return the error description or an empty string if success.

Function used to start the parse algorithm. Creates the structures needed to manage the definitions included in file file.

Returns: ASN1_SUCCESS if the file has a correct syntax and every identifier is known,

ASN1_ELEMENT_NOT_EMPTY if definitions not NULL , ASN1_FILE_NOT_FOUND if an

error occured while opening file , ASN1_SYNTAX_ERROR if the syntax is not correct, ASN1_IDENTIFIER_NOT_FOUND if in the file there is an identifier that is not defined, ASN1_NAME_TOO_LONG if in the file there is an identifier whith more than ASN1_MAX_ NAME_SIZE characters.

asn1 parser2array

[Function] int asn1_parser2array (const char * inputFileName, const char *
   outputFileName, const char * vectorName, char * error_desc)

inputFileName: specify the path and the name of file that contains ASN.1 declarations.

outputFileName: specify the path and the name of file that will contain the C vector definition.

vectorName: specify the name of the C vector.

error desc: return the error description or an empty string if success.

Function that generates a C structure from an ASN1 file. Creates a file containing a C vector to use to manage the definitions included in inputFileName file.

If inputFileName is "/aa/bb/xx.yy" and outputFileName is NULL , the file cre-

ated is "/aa/bb/xx asn1 tab.c". If vectorName is NULL the vector name will be "xx asn1 tab".

Returns: ASN1_SUCCESS if the file has a correct syntax and every identifier is known,

ASN1_FILE_NOT_FOUND if an error occured while opening inputFileName , ASN1_
SYNTAX_ERROR if the syntax is not correct, ASN1_IDENTIFIER_NOT_FOUND if in the file

there is an identifier that is not defined, ASN1_NAME_TOO_LONG if in the file there is an identifier whith more than ASN1_MAX_NAME_SIZE characters.

4.2 ASN.1 field functions

asn1 array2tree

[Function] int asn1_array2tree (const asn1 static node * array, asn1 node *

   definitions, char * errorDescription)

array: specify the array that contains ASN.1 declarations

definitions: return the pointer to the structure created by *ARRAY ASN.1 declarations

errorDescription: return the error description.

Creates the structures needed to manage the ASN.1 definitions. array is a vector

created by asn1_parser2array() .

Returns: ASN1_SUCCESS if structure was created correctly, ASN1_ELEMENT_NOT_EMPTY

if * definitions not NULL, ASN1_IDENTIFIER_NOT_FOUND if in the file there is an

identifier that is not defined (see errorDescription for more information), ASN1_ ARRAY_ERROR if the array pointed by array is wrong.

asn1 delete structure

[Function] int asn1_delete_structure (asn1 node * structure)

structure: pointer to the structure that you want to delete.

Deletes the structure * structure . At the end, * structure is set to NULL.

Returns: ASN1_SUCCESS if successful, ASN1_ELEMENT_NOT_FOUND if * structure was

NULL.

asn1 delete structure2

[Function] int asn1_delete_structure2 (asn1 node * structure, unsigned int
   flags)

structure: pointer to the structure that you want to delete.

flags: additional flags (see ASN1_DELETE_FLAG )

Deletes the structure * structure . At the end, * structure is set to NULL.

Returns: ASN1_SUCCESS if successful, ASN1_ELEMENT_NOT_FOUND if * structure was

NULL.

asn1 delete element

[Function] int asn1_delete_element (asn1 node structure, const char *
   element_name)

structure: pointer to the structure that contains the element you want to delete.

element name: element’s name you want to delete.

Deletes the element named * element_name inside * structure .

Returns: ASN1_SUCCESS if successful, ASN1_ELEMENT_NOT_FOUND if the element_name

was not found.

asn1 create element

[Function] int asn1_create_element (asn1 node definitions, const char *
   source_name, asn1 node * element)

definitions: pointer to the structure returned by "parser asn1" function source name: the name of the type of the new structure (must be inside p structure). element: pointer to the structure created. Creates a structure of type source_name . Example using "pkix.asn": rc = asn1 create element(cert def, "PKIX1.Certificate", certptr);

Returns: ASN1_SUCCESS if creation OK, ASN1_ELEMENT_NOT_FOUND if source_name

is not known.

asn1 print structure

[Function] void asn1_print_structure (FILE * out, asn1 node structure,

const char * name, int mode) out: pointer to the output file (e.g. stdout). structure: pointer to the structure that you want to visit. name: an element of the structure mode: specify how much of the structure to print, can be ASN1_PRINT_NAME , ASN1_

PRINT_NAME_TYPE , ASN1_PRINT_NAME_TYPE_VALUE , or ASN1_PRINT_ALL .

Prints on the out file descriptor the structure’s tree starting from the name element inside the structure structure .

asn1 number of elements

[Function] int asn1_number_of_elements (asn1 node element, const char *
   name, int * num)

element: pointer to the root of an ASN1 structure. name: the name of a sub-structure of ROOT. num: pointer to an integer where the result will be stored Counts the number of elements of a sub-structure called NAME with names equal to "?1","?2", ...

Returns: ASN1_SUCCESS if successful, ASN1_ELEMENT_NOT_FOUND if name is not known,
ASN1_GENERIC_ERROR if pointer num is NULL .

asn1 find structure from oid

[Function] const char * asn1_find_structure_from_oid (asn1 node
   definitions, const char * oidValue)

definitions: ASN1 definitions oidValue: value of the OID to search (e.g. "1.2.3.4"). Search the structure that is defined just after an OID definition. Returns: NULL when oidValue not found, otherwise the pointer to a constant string that contains the element name defined just after the OID.

asn1 copy node

[Function] int asn1_copy_node (asn1 node dst, const char * dst_name,

asn1 node src, const char * src_name) dst: Destination asn1 node. dst name: Field name in destination node. src: Source asn1 node. src name: Field name in source node. Create a deep copy of a asn1 node variable. That function requires dst to be ex-

panded using asn1_create_element() .

Returns: Return ASN1_SUCCESS on success.

asn1 dup node

[Function] asn1_node asn1_dup_node (asn1 node src, const char * src_name)

src: Source asn1 node. src name: Field name in source node. Create a deep copy of a asn1 node variable. This function will return an exact copy of the provided structure. Returns: Return NULL on failure.

asn1 write value

[Function] int asn1_write_value (asn1 node node_root, const char * name,

const void * ivalue, int len) node root: pointer to a structure name: the name of the element inside the structure that you want to set. ivalue: vector used to specify the value to set. If len is >0, VALUE must be a two’s complement form integer. if len=0 *VALUE must be a null terminated string with an integer value. len: number of bytes of *value to use to set the value: value[0]..value[len-1] or 0 if value is a null terminated string Set the value of one element inside a structure. If an element is OPTIONAL and you want to delete it, you must use the value=NULL and len=0. Using "pkix.asn": result=asn1 write value(cert, "tbsCertificate.issuerUniqueID", NULL, 0); Description for each type: INTEGER: VALUE must contain a two’s complement form integer. value[0]=0xFF , len=1 -> integer=-1. value[0]=0xFF value[1]=0xFF , len=2 -> integer=-1. value[0]=0x01 , len=1 -> integer= 1. value[0]=0x00 value[1]=0x01 , len=2 -> integer= 1. value="123" , len=0 -> integer= 123. ENUMERATED: As INTEGER (but only with not negative numbers). BOOLEAN: VALUE must be the null terminated string "TRUE" or "FALSE" and LEN != 0. value="TRUE" , len=1 -> boolean=TRUE. value="FALSE" , len=1 -> boolean=FALSE. OBJECT IDENTIFIER: VALUE must be a null terminated string with each number separated by a dot (e.g. "1.2.3.543.1"). LEN != 0. value="1 2 840 10040 4 3" , len=1 -> OID=dsa-with-sha. UTCTime: VALUE must be a null terminated string in one of these formats: "YYMMDDhhmmssZ", "YYMMDDhhmmssZ", "YYMMDDhhmmss+hh’mm’", "YYMMDDhhmmss-hh’mm’", "YYMMDDhhmm+hh’mm’", or "YYMMDDhhmmhh’mm’". LEN != 0. value="9801011200Z" , len=1 -> time=Jannuary 1st, 1998 at 12h 00m Greenwich Mean Time GeneralizedTime: VALUE must be in one of this format: "YYYYMMDDhh-

mmss.sZ",  "YYYYMMDDhhmmss.sZ",  "YYYYMMDDhhmmss.s+hh’mm’",
"YYYYMMDDhhmmss.s-hh’mm’",  "YYYYMMDDhhmm+hh’mm’",  or

"YYYYMMDDhhmm-hh’mm’" where ss.s indicates the seconds with any precision like "10.1" or "01.02". LEN != 0 value="2001010112001.12-0700" , len=1 -> time=Jannuary 1st, 2001 at 12h 00m 01.12s Pacific Daylight Time OCTET STRING: VALUE contains the octet string and LEN is the number of octets. value="$\backslash$x01$\backslash$x02$\backslash$x03" , len=3 -> three bytes octet string GeneralString: VALUE contains the generalstring and LEN is the number of octets. value="$\backslash$x01$\backslash$x02$\backslash$x03" , len=3 -> three bytes generalstring BIT STRING: VALUE contains the bit string organized by bytes and LEN is the number of bits. value="$\backslash$xCF" , len=6 -> bit string="110011" (six bits) CHOICE: if NAME indicates a choice type, VALUE must specify one of the alternatives with a null terminated string. LEN != 0. Using "pkix.asn"\: result=asn1 write value(cert, "certificate1.tbsCertificate.subject", "rdnSequence", 1); ANY: VALUE indicates the der encoding of a structure. LEN != 0. SEQUENCE OF: VALUE must be the null terminated string "NEW" and LEN != 0. With this instruction another element is appended in the sequence. The name of this element will be "?1" if it’s the first one, "?2" for the second and so on. Using "pkix.asn"\: result=asn1 write value(cert, "certificate1.tbsCertificate.subject.rdnSequence", "NEW", 1); SET OF: the same as SEQUENCE OF. Using "pkix.asn": result=asn1 write value(cert, "tbsCertificate.subject.rdnSequence.?LAST", "NEW", 1);

Returns: ASN1_SUCCESS if the value was set, ASN1_ELEMENT_NOT_FOUND if name is not

a valid element, and ASN1_VALUE_NOT_VALID if ivalue has a wrong format.

asn1 read value

[Function] int asn1_read_value (asn1 node root, const char * name, void *

   ivalue, int * len)

root: pointer to a structure. name: the name of the element inside a structure that you want to read. ivalue: vector that will contain the element’s content, must be a pointer to memory cells already allocated (may be NULL ). len: number of bytes of *value: value[0]..value[len-1]. Initialy holds the sizeof value. Returns the value of one element inside a structure. If an element is OPTIONAL and this returns ASN1_ELEMENT_NOT_FOUND , it means that this element wasn’t present in the der encoding that created the structure. The first element of a SEQUENCE OF or SET OF is named "?1". The second one "?2" and so on. If the root provided is a node to specific sequence element, then the keyword "?CURRENT" is also acceptable and indicates the current sequence element of this node. Note that there can be valid values with length zero. In these case this function will succeed and len will be zero. INTEGER: VALUE will contain a two’s complement form integer. integer=-1 -> value[0]=0xFF , len=1. integer=1 -> value[0]=0x01 , len=1. ENUMERATED: As INTEGER (but only with not negative numbers). BOOLEAN: VALUE will be the null terminated string "TRUE" or "FALSE" and LEN=5 or LEN=6. OBJECT IDENTIFIER: VALUE will be a null terminated string with each number separated by a dot (i.e. "1.2.3.543.1"). LEN = strlen(VALUE)+1 UTCTime: VALUE will be a null terminated string in one of these formats: "YYM-MDDhhmmss+hh’mm’" or "YYMMDDhhmmss-hh’mm’". LEN=strlen(VALUE)+1. GeneralizedTime: VALUE will be a null terminated string in the same format used to set the value. OCTET STRING: VALUE will contain the octet string and LEN will be the number of octets. GeneralString: VALUE will contain the generalstring and LEN will be the number of octets. BIT STRING: VALUE will contain the bit string organized by bytes and LEN will be the number of bits. CHOICE: If NAME indicates a choice type, VALUE will specify the alternative selected. ANY: If NAME indicates an any type, VALUE will indicate the DER encoding of the structure actually used.

Returns: ASN1_SUCCESS if value is returned, ASN1_ELEMENT_NOT_FOUND if name is

not a valid element, ASN1_VALUE_NOT_FOUND if there isn’t any value for the element selected, and ASN1_MEM_ERROR if The value vector isn’t big enough to store the result, and in this case len will contain the number of bytes needed.

asn1 read value type

[Function] int asn1_read_value_type (asn1 node root, const char * name, void *

ivalue, int * len, unsigned int * etype) root: pointer to a structure. name: the name of the element inside a structure that you want to read. ivalue: vector that will contain the element’s content, must be a pointer to memory cells already allocated (may be NULL ). len: number of bytes of *value: value[0]..value[len-1]. Initialy holds the sizeof value. etype: The type of the value read (ASN1 ETYPE) Returns the type and value of one element inside a structure. If an element is OP-TIONAL and this returns ASN1_ELEMENT_NOT_FOUND , it means that this element wasn’t present in the der encoding that created the structure. The first element of a SEQUENCE OF or SET OF is named "?1". The second one "?2" and so on. If the root provided is a node to specific sequence element, then the keyword "?CUR-RENT" is also acceptable and indicates the current sequence element of this node. Note that there can be valid values with length zero. In these case this function will succeed and len will be zero. INTEGER: VALUE will contain a two’s complement form integer. integer=-1 -> value[0]=0xFF , len=1. integer=1 -> value[0]=0x01 , len=1. ENUMERATED: As INTEGER (but only with not negative numbers). BOOLEAN: VALUE will be the null terminated string "TRUE" or "FALSE" and LEN=5 or LEN=6. OBJECT IDENTIFIER: VALUE will be a null terminated string with each number separated by a dot (i.e. "1.2.3.543.1"). LEN = strlen(VALUE)+1 UTCTime: VALUE will be a null terminated string in one of these formats: "YYM-MDDhhmmss+hh’mm’" or "YYMMDDhhmmss-hh’mm’". LEN=strlen(VALUE)+1. GeneralizedTime: VALUE will be a null terminated string in the same format used to set the value. OCTET STRING: VALUE will contain the octet string and LEN will be the number of octets. GeneralString: VALUE will contain the generalstring and LEN will be the number of octets. BIT STRING: VALUE will contain the bit string organized by bytes and LEN will be the number of bits. CHOICE: If NAME indicates a choice type, VALUE will specify the alternative selected. ANY: If NAME indicates an any type, VALUE will indicate the DER encoding of the structure actually used.

Returns: ASN1_SUCCESS if value is returned, ASN1_ELEMENT_NOT_FOUND if name is

not a valid element, ASN1_VALUE_NOT_FOUND if there isn’t any value for the element selected, and ASN1_MEM_ERROR if The value vector isn’t big enough to store the result, and in this case len will contain the number of bytes needed.

asn1 read tag

[Function] int asn1_read_tag (asn1 node root, const char * name, int *

   tagValue, int * classValue)

root: pointer to a structure name: the name of the element inside a structure. tagValue: variable that will contain the TAG value. classValue: variable that will specify the TAG type. Returns the TAG and the CLASS of one element inside a structure. CLASS can

have one of these constants: ASN1_CLASS_APPLICATION , ASN1_CLASS_UNIVERSAL ,
ASN1_CLASS_PRIVATE or ASN1_CLASS_CONTEXT_SPECIFIC .
Returns: ASN1_SUCCESS if successful, ASN1_ELEMENT_NOT_FOUND if name is not a valid

element.

asn1 read node value

[Function] int asn1_read_node_value (asn1 node node, asn1 data node st *

   data)

node: pointer to a node. data: a point to a asn1 data node st Returns the value a data node inside a asn1 node structure. The data returned should be handled as constant values. Returns: ASN1_SUCCESS if the node exists.

4.3 DER functions

asn1 length der

[Function] void asn1_length_der (unsigned long int len, unsigned char * der, int

   * der_len)

len: value to convert. der: buffer to hold the returned encoding (may be NULL ). der len: number of meaningful bytes of ANS (der[0]..der[der len-1]). Creates the DER encoding of the provided length value. The der buffer must have enough room for the output. The maximum length this function will encode is ASN1_

MAX_LENGTH_SIZE .

To know the size of the DER encoding use a NULL value for der .

asn1 octet der

[Function] void asn1_octet_der (const unsigned char * str, int str_len,

unsigned char * der, int * der_len) str: the input data. str len: STR length (str[0]..str[*str len-1]). der: encoded string returned. der len: number of meaningful bytes of DER (der[0]..der[der len-1]). Creates a length-value DER encoding for the input data. The DER encoding of the input data will be placed in the der variable. Note that the OCTET STRING tag is not included in the output. This function does not return any value because it is expected that der_len will contain enough bytes to store the string plus the DER encoding. The DER encoding size can be obtained using asn1_length_der() .

asn1 encode simple der

[Function] int asn1_encode_simple_der (unsigned int etype, const unsigned

char * str, unsigned int str_len, unsigned char * tl, unsigned int * tl_len) etype: The type of the string to be encoded (ASN1 ETYPE ) str: the string data. str len: the string length tl: the encoded tag and length tl len: the bytes of the tl field Creates the DER encoding for various simple ASN.1 types like strings etc. It stores the tag and length in tl , which should have space for at least ASN1_MAX_TL_SIZE bytes. Initially tl_len should contain the size of tl . The complete DER encoding should consist of the value in tl appended with the provided str . Returns: ASN1_SUCCESS if successful or an error value.

asn1 bit der

[Function] void asn1_bit_der (const unsigned char * str, int bit_len, unsigned

char * der, int * der_len) str: BIT string. bit len: number of meaningful bits in STR. der: string returned. der len: number of meaningful bytes of DER (der[0]..der[ans len-1]). Creates a length-value DER encoding for the input data as it would have been for a BIT STRING. The DER encoded data will be copied in der . Note that the BIT STRING tag is not included in the output. This function does not return any value because it is expected that der_len will contain enough bytes to store the string plus the DER encoding. The DER encoding size can be obtained using asn1_length_der() .

asn1 der coding

[Function] int asn1_der_coding (asn1 node element, const char * name, void *

   ider, int * len, char * ErrorDescription)

element: pointer to an ASN1 element name: the name of the structure you want to encode (it must be inside *POINTER). ider: vector that will contain the DER encoding. DER must be a pointer to memory cells already allocated. len: number of bytes of * ider : ider [0].. ider [len-1], Initialy holds the sizeof of der vector. ErrorDescription: return the error description or an empty string if success. Creates the DER encoding for the NAME structure (inside *POINTER structure).

Returns: ASN1_SUCCESS if DER encoding OK, ASN1_ELEMENT_NOT_FOUND if name is

not a valid element, ASN1_VALUE_NOT_FOUND if there is an element without a value, ASN1_MEM_ERROR if the ider vector isn’t big enough and in this case len will contain the length needed.

asn1 get length der

[Function] long asn1_get_length_der (const unsigned char * der, int der_len,

int * len) der: DER data to decode. der len: Length of DER data to decode. len: Output variable containing the length of the DER length field. Extract a length field from DER data. Returns: Return the decoded length value, or -1 on indefinite length, or -2 when the value was too big to fit in a int, or -4 when the decoded length value plus len would

exceed der_len .

asn1 get tag der

[Function] int asn1_get_tag_der (const unsigned char * der, int der_len,

unsigned char * cls, int * len, unsigned long * tag) der: DER data to decode. der len: Length of DER data to decode. cls: Output variable containing decoded class. len: Output variable containing the length of the DER TAG data. tag: Output variable containing the decoded tag. Decode the class and TAG from DER code. Returns: Returns ASN1_SUCCESS on success, or an error.

asn1 get length ber

[Function] long asn1_get_length_ber (const unsigned char * ber, int ber_len,

int * len) ber: BER data to decode. ber len: Length of BER data to decode. len: Output variable containing the length of the BER length field.

Extract a length field from BER data. The difference to asn1_get_length_der() is that this function will return a length even if the value has indefinite encoding. Returns: Return the decoded length value, or negative value when the value was too big. Since: 2.0

asn1 get octet der

[Function] int asn1_get_octet_der (const unsigned char * der, int der_len, int

   * ret_len, unsigned char * str, int str_size, int * str_len)

der: DER data to decode containing the OCTET SEQUENCE. der len: Length of DER data to decode. ret len: Output variable containing the length of the DER data. str: Pre-allocated output buffer to put decoded OCTET SEQUENCE in. str size: Length of pre-allocated output buffer. str len: Output variable containing the length of the OCTET SEQUENCE. Extract an OCTET SEQUENCE from DER data. Returns: Returns ASN1_SUCCESS on success, or an error.

asn1 get bit der

[Function] int asn1_get_bit_der (const unsigned char * der, int der_len, int *

   ret_len, unsigned char * str, int str_size, int * bit_len)

der: DER data to decode containing the BIT SEQUENCE. der len: Length of DER data to decode. ret len: Output variable containing the length of the DER data. str: Pre-allocated output buffer to put decoded BIT SEQUENCE in. str size: Length of pre-allocated output buffer. bit len: Output variable containing the size of the BIT SEQUENCE. Extract a BIT SEQUENCE from DER data. Returns: Return ASN1_SUCCESS on success, or an error.

asn1 der decoding2

[Function] int asn1_der_decoding2 (asn1 node * element, const void * ider, int

   * max_ider_len, unsigned int flags, char * errorDescription)

element: pointer to an ASN1 structure. ider: vector that contains the DER encoding. max ider len: pointer to an integer giving the information about the maximal number of bytes occupied by * ider . The real size of the DER encoding is returned through this pointer. flags: flags controlling the behaviour of the function. errorDescription: null-terminated string contains details when an error occurred.

Fill the structure * element with values of a DER encoding string. The structure must just be created with function asn1_create_element() .

If ASN1_DECODE_FLAG_ALLOW_PADDING flag is set then the function will ignore padding after the decoded DER data. Upon a successful return the value of * max_ider_len will be set to the number of bytes decoded.

Returns: ASN1_SUCCESS if DER encoding OK, ASN1_ELEMENT_NOT_FOUND if ELEMENT

is NULL , and ASN1_TAG_ERROR or ASN1_DER_ERROR if the der encoding doesn’t match the structure name (* ELEMENT deleted).

asn1 der decoding

[Function] int asn1_der_decoding (asn1 node * element, const void * ider, int

   ider_len, char * errorDescription)

element: pointer to an ASN1 structure.

ider: vector that contains the DER encoding.

ider len: number of bytes of * ider : ider [0].. ider [len-1].

errorDescription: null-terminated string contains details when an error occurred.

Fill the structure * element with values of a DER encoding string. The structure must just be created with function asn1_create_element() .

Note that the * element variable is provided as a pointer for historical reasons.

Returns: ASN1_SUCCESS if DER encoding OK, ASN1_ELEMENT_NOT_FOUND if ELEMENT

is NULL , and ASN1_TAG_ERROR or ASN1_DER_ERROR if the der encoding doesn’t match the structure name (* ELEMENT deleted).

asn1 der decoding element

[Function] int asn1_der_decoding_element (asn1 node * structure, const char
   * elementName, const void * ider, int len, char * errorDescription)

structure: pointer to an ASN1 structure

elementName: name of the element to fill

ider: vector that contains the DER encoding of the whole structure.

len: number of bytes of *der: der[0]..der[len-1]

errorDescription: null-terminated string contains details when an error occurred.

Fill the element named ELEMENTNAME with values of a DER encoding string. The structure must just be created with function asn1_create_element() . The DER vector must contain the encoding string of the whole STRUCTURE . If an error occurs during the decoding procedure, the * STRUCTURE is deleted and set equal to NULL .

This function is deprecated and may just be an alias to asn1 der decoding in future versions. Use asn1_der_decoding() instead.

Returns: ASN1_SUCCESS if DER encoding OK, ASN1_ELEMENT_NOT_FOUND if ELE-
MENT is NULL or elementName == NULL, and ASN1_TAG_ERROR or ASN1_DER_ERROR

if the der encoding doesn’t match the structure structure (*ELEMENT deleted).

asn1 der decoding startEnd

[Function] int asn1_der_decoding_startEnd (asn1 node element, const void *
   ider, int ider_len, const char * name_element, int * start, int * end)

element: pointer to an ASN1 element ider: vector that contains the DER encoding. ider len: number of bytes of * ider : ider [0].. ider [len-1] name element: an element of NAME structure. start: the position of the first byte of NAME ELEMENT decoding ( ider [*start]) end: the position of the last byte of NAME ELEMENT decoding ( ider [*end]) Find the start and end point of an element in a DER encoding string. I mean that if you have a der encoding and you have already used the function asn1_der_ decoding() to fill a structure, it may happen that you want to find the piece of string concerning an element of the structure. One example is the sequence "tbsCertificate" inside an X509 certificate. Note that since libtasn1 3.7 the ider and ider_len parameters can be omitted, if the element is already decoded using asn1_der_decoding() .

Returns: ASN1_SUCCESS if DER encoding OK, ASN1_ELEMENT_NOT_FOUND if ELE-

MENT is asn1_node EMPTY or name_element is not a valid element, ASN1_TAG_ ERROR or ASN1_DER_ERROR if the der encoding doesn’t match the structure ELE-MENT.

asn1 expand any defined by

[Function] int asn1_expand_any_defined_by (asn1 node definitions,

asn1 node * element) definitions: ASN1 definitions element: pointer to an ASN1 structure Expands every "ANY DEFINED BY" element of a structure created from a DER decoding process (asn1 der decoding function). The element ANY must be defined by an OBJECT IDENTIFIER. The type used to expand the element ANY is the first one following the definition of the actual value of the OBJECT IDENTIFIER. Returns: ASN1_SUCCESS if Substitution OK, ASN1_ERROR_TYPE_ANY if some "ANY DEFINED BY" element couldn’t be expanded due to a problem in OBJECT ID -> TYPE association, or other error codes depending on DER decoding.

asn1 expand octet string

[Function] int asn1_expand_octet_string (asn1 node definitions, asn1 node
   * element, const char * octetName, const char * objectName)

definitions: ASN1 definitions element: pointer to an ASN1 structure octetName: name of the OCTECT STRING field to expand. objectName: name of the OBJECT IDENTIFIER field to use to define the type for expansion.

Expands an "OCTET STRING" element of a structure created from a DER decoding process (the asn1_der_decoding() function). The type used for expansion is the first one following the definition of the actual value of the OBJECT IDENTIFIER indicated by OBJECTNAME.

Returns: ASN1_SUCCESS if substitution OK, ASN1_ELEMENT_NOT_FOUND if objectName

or octetName are not correct, ASN1_VALUE_NOT_VALID if it wasn’t possible to find the type to use for expansion, or other errors depending on DER decoding.

asn1 decode simple der

[Function] int asn1_decode_simple_der (unsigned int etype, const unsigned

char * der, unsigned int der_len, const unsigned char ** str, unsigned int *

   str_len)

etype: The type of the string to be encoded (ASN1 ETYPE )

der: the encoded string

der len: the bytes of the encoded string

str: a pointer to the data

str len: the length of the data

Decodes a simple DER encoded type (e.g. a string, which is not constructed). The output is a pointer inside the der .

Returns: ASN1_SUCCESS if successful or an error value.

4.4 Error handling functions

asn1 perror

[Function] void asn1_perror (int error)

error: is an error returned by a libtasn1 function.

Prints a string to stderr with a description of an error. This function is like perror() . The only difference is that it accepts an error returned by a libtasn1 function.

Since: 1.6

asn1 strerror

[Function] const char * asn1_strerror (int error)

error: is an error returned by a libtasn1 function.

Returns a string with a description of an error. This function is similar to strerror. The only difference is that it accepts an error (number) returned by a libtasn1 function.

Returns: Pointer to static zero-terminated string describing error code.

Since: 1.6

4.5 Auxilliary functions

asn1 find node

[Function] asn1_node asn1_find_node (asn1 node pointer, const char * name)

pointer: NODE ASN element pointer. name: null terminated string with the element’s name to find. Searches for an element called name starting from pointer . The name is composed by differents identifiers separated by dots. When * pointer has a name, the first identifier must be the name of * pointer , otherwise it must be the name of one child

of * pointer .

Returns: the search result, or NULL if not found.

asn1 check version

[Function] const char * asn1_check_version (const char * req_version)

req version: Required version number, or NULL . Check that the version of the library is at minimum the requested one and return the version string; return NULL if the condition is not satisfied. If a NULL is passed to this function, no check is done, but the version string is simply returned.

See ASN1_VERSION for a suitable req_version string.

Returns: Version string of run-time library, or NULL if the run-time library does not meet the required version number.

Appendix A Copying Information

A.1 GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING

You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public.

It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS

You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version:

A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 30

Concept Index

A

ASN.1 schema . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 asn1Coding program . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 asn1Decoding program. . . . . . . . . . . . . . . . . . . . . . . . . . 6 asn1Parser program. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

FDL, GNU Free Documentation License . . . . . . . 22 Future developments. . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

H

Header file libtasn1.h . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

M

Main type asn1 node . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

P

Porting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

S

Supported ASN.1 types, list of . . . . . . . . . . . . . . . . . . 2

T

threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Function and Data Index 31

Function and Data Index

asn1_array2tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 asn1_bit_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 asn1_check_version . . . . . . . . . . . . . . . . . . . . . . . . . . 21 asn1_copy_node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 asn1_create_element . . . . . . . . . . . . . . . . . . . . . . . . . . 9 asn1_decode_simple_der . . . . . . . . . . . . . . . . . . . . . . 20 asn1_delete_element . . . . . . . . . . . . . . . . . . . . . . . . . . 8 asn1_delete_structure . . . . . . . . . . . . . . . . . . . . . . . . 8 asn1_delete_structure2 . . . . . . . . . . . . . . . . . . . . . . . 8 asn1_der_coding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 asn1_der_decoding. . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

asn1_der_decoding_element . . . . . . . . . . . . . . . . . . 18
asn1_der_decoding_startEnd . . . . . . . . . . . . . . . . . 19

asn1_der_decoding2 . . . . . . . . . . . . . . . . . . . . . . . . . . 17 asn1_dup_node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 asn1_encode_simple_der . . . . . . . . . . . . . . . . . . . . . . 15

asn1_expand_any_defined_by . . . . . . . . . . . . . . . . . 19
asn1_expand_octet_string . . . . . . . . . . . . . . . . . . . 19

asn1_find_node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

                                                asn1_find_structure_from_oid . . . . . . . . . . . . . . . . 9

asn1_get_bit_der. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17
asn1_get_length_ber . . . . . . . . . . . . . . . . . . . . . . . . . 16
asn1_get_length_der . . . . . . . . . . . . . . . . . . . . . . . . . 16
asn1_get_octet_der . . . . . . . . . . . . . . . . . . . . . . . . . . 17
asn1_get_tag_der. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
asn1_length_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

                                                asn1_number_of_elements . . . . . . . . . . . . . . . . . . . . . . 9

asn1_octet_der . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14
asn1_parser2array. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
asn1_parser2tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
asn1_perror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
asn1_print_structure . . . . . . . . . . . . . . . . . . . . . . . . . 9
asn1_read_node_value . . . . . . . . . . . . . . . . . . . . . . . . 14
asn1_read_tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14
asn1_read_value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12
asn1_read_value_type . . . . . . . . . . . . . . . . . . . . . . . . 13
asn1_strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
asn1_write_value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10